OpenStudy (anonymous):

Jake's band played songs at three different parties. At the first party they played four-fifths as long as they did at the second party, which was one-half as long as the third party. If they played for 6 minutes at the first party, how many minutes did they play at the third party?

OpenStudy (anonymous):

@dumbcow pls help me

OpenStudy (dumbcow):

ok so let x,y,z represent 1st party,2nd party,3rd party x=6 also you are told x = (4/5)y y = (1/2)z make substitutions 6 = 4/5 y y = 30/4 = 7.5 7.5 = 1/2 z z = 15

OpenStudy (anonymous):

@dumbcow ty so much and ty for the explanation

OpenStudy (dumbcow):

yw